### 09 Apr 2024 v1

# Wu’s Method can Boost Symbolic AI to Rival Silver Medalists and AlphaGeometry to Outperform Gold Medalists at IMO Geometry

Abstract
Proving geometric theorems constitutes a hallmark of visual reasoning combining both intuitive and logical skills. Therefore, automated theorem proving of Olympiad-level geometry problems is considered a notable milestone in human-level automated reasoning. The introduction of AlphaGeometry, a neuro-symbolic model trained with 100 million synthetic samples, marked a major breakthrough. It solved 25 of 30 International Mathematical Olympiad (IMO) problems whereas the reported baseline based on Wu’s method solved only ten. In this note, we revisit the IMO-AG-30 Challenge introduced with AlphaGeometry, and find that Wu’s method is surprisingly strong. Wu’s method alone can solve 15 problems, and some of them are not solved by any of the other methods. This leads to two key findings: (i) Combining Wu’s method with the classic synthetic methods of deductive databases and angle, ratio, and distance chasing solves 21 out of 30 methods by just using a CPU-only laptop with a time limit of 5 minutes per problem. Essentially, this classic method solves just 4 problems less than AlphaGeometry and establishes the first fully symbolic baseline, strong enough to rival the performance of an IMO silver medalist. (ii) Wu’s method even solves 2 of the 5 problems that AlphaGeometry failed to solve. Thus, by combining AlphaGeometry with Wu’s method we set a new state-of-the-art for automated theorem proving on IMO-AG-30, solving 27 out of 30 problems, the first AI method which outperforms an IMO gold medalist.
[…]
1 Introduction
Automated theorem proving has been the long-term goal of developing computer programs that can match the conjecturing and proving capabilities demanded by mathematical research [10]. This field has recognized solving Olympiad-level geometry problems as a key milestone [2, 1], marking a frontier of computers to perform complex mathematical reasoning. The International Mathematical Olympiad (IMO) started in 1959 and hosts the most reputed theorem-proving competitions in the world that play an important role in identifying exceptional talents in problem solving. In fact, half of all Fields medalists participated in the IMO in their youth, and matching top human performances at the olympiad level has become a notable milestone of AI research.
Euclidean geometry is well suited to testing the reasoning skills of AI systems. It is finitely axiomatized [15] and many proof systems for Euclidean geometry have been proposed over the years which are amenable to automated theorem proving techniques [4, 5]. Furthermore, proof search can be guided by diagrammatic representations [12, 18], probabilistic verification [11, 23], and a vast array of possible deductions using human-designed heuristics for properties like angles, areas, and distances, methods affectionately called “trig bashing” and “bary bashing[24, 25] by International Mathematical Olympiad (IMO) participants. In addition, this domain is challenging — specific proof systems need to be defined for specifying the problem, there is a shortage of data to train from, and problems typically contain ambiguities around degenerate cases [29, 20, 17] that are complex to resolve and have led to the humorous folklore that “geometry problems never take care of degeneracies”.
Automated reasoning in geometry can be categorized into algebraic [28, 27, 16] and synthetic methods [12, 6, 22]. Recent focus has been on synthetic methods like Deductive Databases (DD) [6] that mimic human-like proving techniques and produce intelligible proofs by treating the problem of theorem proving as a step-by-step search problem using a set of geometry axioms. For instance, DD uses a fixed set of expert-curated geometric rules which are applied repeatedly to an initial geometric configuration. This is performed until the system reaches a fixpoint and no new facts can be deduced using the available rules. AlphaGeometry [26], a novel neuro-symbolic prover, represents a recent breakthrough advancement in this area. It adds additional rules to the prior work of DD to perform angle, ratio, and distance chasing (AR), and the resulting symbolic engine (DD+AR) is further enhanced using constructions suggested by a large language model (DD+AR+LLM-Constructions) trained on 100 million synthetic proofs. It has outclassed algebraic methods by solving 25 of 30 IMO problems, whereas the reported baseline based on Wu’s method [28, 8] solved only ten [26].
Algebraic methods, such as Wu’s method and the Gröbner basis method [16], transform geometric hypotheses into system of polynomials to verify conclusions. They offer powerful procedures that are proven to decide statements in broad classes of geometry [8, 16]. More precisely, Wu’s method possesses the capability to address any problem for which the hypotheses and conclusion can be expressed using algebraic equations [7], while simultaneously generating non-degeneracy conditions automatically [29, 17]. This remarkable feature implies that Wu’s method can handle problems not only in plane geometry but also in solid and higher-dimensional geometries, i.e. in areas where synthetic methods can be used only with great effort and additional considerations. [9].
In this work, we put the capabilities of Wu’s method to the test on problems that involve meaningful and structured mathematical statements, rather than often arbitrary worst case instances [13]. To this end, we re-evaluate Wu’s Method on the IMO-AG-30 benchmark introduced by Trinh et al. [26]. We find that it performs surprisingly strong, solving 15 problems, and some of them are not solved by any of the other methods. This leads to two key findings:
  • Combining Wu’s method (Wu) with the classic synthetic methods of deductive databases (DD) and angle, ratio, and distance chasing (AR) solves 21 out of 30 methods by just using a CPU-only laptop with a time limit of 5 minutes per problem. Essentially, this classic method (Wu&DD+AR) solves just 4 problems less than AlphaGeometry and establishes the first fully symbolic baseline, strong enough to rival the performance of an IMO silver medalist.
  • Wu’s method even solves 2 of the 5 problems that AlphaGeometry (AG) failed to solve. Thus, by combining AlphaGeometry with Wu’s method (Wu&AG) we set a new state-of-the-art for automated theorem proving on IMO-AG-30, solving 27 out of 30 problems, the first AI method which outperforms an IMO gold medalist.
2 Experiments & Results
Figure 1: A) Performance across symbolic and LLM-Augmented methods on the IMO-AG-30 problem set, along with human performance. We set a strong baseline among symbolic systems at the standard of a silver medalist and outperform a gold medalist by a margin of one problem. Right: Diagram showing how the different methods overlap or complement each other on the IMO-AG-30 problems.
In January 2024, IMO-AG-30 was introduced as a new benchmark by Trinh et al. [26] to demonstrate the skill level of AlphaGeometry. IMO-AG-30 is based on geometry problems collected from the IMO competitions since 2000 and adapted to a narrower, specialized environment for classical geometry used in interactive graphical proof assistants, resulting in a test set of 30 classical geometry problems. The number of problems solved in this benchmark are related to the number of problems solved on average by IMO contestants. As indicated by the gray horizontal lines in Figure 1 (A), bronze, silver and gold medalists on average solved 19.3, 22.9 and 25.9 of these problems, and 15.2 represents the average over all contestants. The specific set of problems that have been collected for IMO-AG-30 are listed in the left column of the diagram in Figure 1 (B).
Experimental Details. We evaluated performance using the IMO-AG-30 benchmark, with baselines and dataset all adopted from Trinh et al. [26]. We only re-implemented Wu’s Method through the JGEX software [19, 30] by manual translation of the IMO-AG-30 problems into JGEX-compatible format [^1] However, 4 out of 30 problems were untranslatable due to lack of appropriate constructions within the JGEX framework, hence our reported is out of 26 problems.. We also successfully reproduced the DD+AR baseline, necessary for our final proposed method from the AlphaGeometry codebase. We manually verified that the hypothesis and conclusion equations generated by JGEX for several problems translated by us were indeed correct.
Results. Our findings, are displayed in Figure 1 in combination with previous results from [26]. Figure 1 (A) compares the number of problems solved, and (B) shows which problems are solved by which method to visualize how the different methods overlap or complement each other. In Figure 1 (A), the performance levels of IMO contestants are indicated by gray horizontal lines, showing gold, silver, bronze, average, and honorable mention level. The performance levels of synthetic symbolic methods are displayed with blue bars and of LLM-augmented neurosymbolic methods are shown with green bars. Our own results obtained with Wu’s method fall into the category of algebraic synthetic methods shown with orange bars. All results for synthetic symbolic methods (blue) or neurosymbolic LLM-augmented methods (green) are adopted from [26].
Our combination of Wu’s method with DD+AR sets a new symbolic baseline (Wu&DD+AR) that outperforms all traditional methods by a margin of 6 problems. It solves 21 of the IMO-AG-30 problems, matching the level of AlphaGeometry without fine-tuning (FT-9M only) shown in the Appendix (Figure 2). Wu’s method achieves this performance with remarkably low computational requirements. On a laptop equipped with an AMD Ryzen 7 5800H processor and 16 GB of RAM, we were able to solve 14 out of 15 problems within 5 seconds and one problem (2015 P4) required 3 minutes. In our experiments, Wu’s method either solves problems almost immediately or the laptop runs out of memory within 5 minutes. Remarkably, two of the fifteen problems we were able to solve with Wu’s method (2021 P3, 2008 P1B) were among the five problems that were too difficult to solve for AlphaGeometry. Thus, by simple ensemble combination between Wu’s method and AlphaGeometry, we obtain the new state-of-the-art solving 27 out of 30 problems on the IMO-AG-30 benchmark as visualized by the green/orange bar (Wu&AG) Figure 1.
3 Conclusions
Overall, our note highlights the potential of algebraic methods in automated geometric reasoning for solving International Mathematical Olympiad (IMO) geometry problems [^2] Peter Novotný similarly proved 11 of the 17 IMO Geometry problems from 1984–2003 using the Gröbner basis method, although only after manually adding non-degeneracy conditions [3] as referenced here., raising the number of problems solved with Wu’s method on IMO-AG-30 from ten to fifteen. Among those fifteen problems are several that are difficult for synthetic methods and their LLM-augmented versions that are currently most popular.
To the best of our knowledge, our symbolic baseline is the only symbolic baseline performing above the average IMO contestant and approaching the performance of an IMO silver medalist on geometry. Similarly, our combination of AlphaGeomtery with Wu’s method is the first AI system to outperform a human gold-medalist at IMO geometry problems. This achievement illustrates the complementarity of algebraic and synthetic methods in this area (see Figure 1 B). The usefulness of algebraic approaches is most obvious from the two problems 2008 P1B and 2021 P3 which are currently solved by no automatic theorem prover other than Wu’s method.
While algebraic methods have always been recognized for their theoretical guarantees, their usefulness has been previously questioned for being too slow and not human interpretable. Our observations indicate that on several problems Wu’s Method performs more efficiently than previously recognized, and we advocate against dismissing it solely on the basis of its inability to produce human-readable proofs.
Our results are work-in-progress, currently hindered by the scarce availability of existing implementations, each with their significant inadequacies including limited constructions and suboptimal performance, despite the theoretical promise. We believe it might be feasible to outperform AlphaGeometry’s proving capabilities through purely traditional methods and hope our note encourages improving current software for classical computational approaches in this area.
On the other hand, the notable success of our approach suggests that IMO geometry problems, while challenging for humans, might not fully test the limits of modern computational solvers. Repeated application of human-defined heuristics combined with few constructions (up to three construction have been previously shown to be sufficient [26]) for achieving success rather than on a deep, complex exploration of combinatorial possibilities. This situation is reminiscent of chess endgames, which were mastered relatively early on by brute-force solvers [14, 21]. We hope this note encourages development of new benchmarks for automated theorem provers in geometry.
IMO Geometry feels like a chess endgame, fixed with little wiggle room.
The real fun lies in teaching machines to solve the Geometry middlegames!
Siddharth Bhat
Acknowledgements
The authors would like to thank (in alphabetic order): Shashwat Goel, Shyamgopal Karthik, Yash Sharma, Matthias Tangemann, Saujas Vaduguru for helpful feedback on the draft. MB acknowledges financial support via the Open Philanthropy Foundation funded by the Good Ventures Foundation.
Appendix A Detailed Comparisons
Extended Comparisons. We compare with all human and automated methods on the IMO-AG-30 benchmark [26] in Figure 2. Our evaluation includes GPT4, Full-Angle method (FA), Gröbner Basis (Gröbner), Deductive Databases (DD), Deductive Databases combined with Algebraic Rules and enhancements with GPT-4 for construction suggestions (DD+AR+GPT4). Additionally, we examined different configurations of the AlphaGeometry model: one only pretrained on 100 million samples (PT-100M) and other only finetuned on 9 million constructions (FT-9M). Note that we construct the Wu&DD+AR baseline by simply parallely running both Wu’s and DD+AR methods and stopping when either method solves the problem. Similarly, we construct the Wu&AlphaGeometry baseline. We see that our Wu&DD+AR baseline matches AG (FT-9M) baseline while Wu’s method alone matches the best DD+AR+GPT4 algorithm.
Figure 2: Left: Performance across symbolic and LLM-Augmented methods on the IMO-AG-30 problem set, along with human performance. We set a strong baseline among symbolic systems at the standard of a silver medalist and outperform a gold medalist by a margin of one problem. Right: Diagram showing how the different methods overlap or complement each other on the IMO-AG-30 problems.
Appendix B Illustrations: 2008 P1B and 2021 P3
We provide illustrations of the solutions of Wu’s method for the two problems AlphaGeometry could not solve to allow for additional scrutiny without having to reproduce the same on the JGEX solver.
Figure 3: Problem 2008-P1B JGEX (Above) and 2021-P3 (Below) with Input (Left) and Generated Solution (Right) for Wu’s method. This illustration can be reproduced by opening the .gex files provided alongside on the HuggingFace repository and pressing Run.